OpenStudy (anonymous):

What is the value of x? 1/4x-8+2/4x=19

OpenStudy (anonymous):

36

OpenStudy (anonymous):

1/4x - 8 + 2/4x = 19 Add 8 to both sides 1/4x + 2/4x = 27 Combine like terms 3/4x = 27 Multiply 3/4 to both sides x = 20.25 I think I messed up on this one so I may need someone to check.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I have a CC to attend. I won't be able to help you this next hour. Sorry.

OpenStudy (anonymous):

ok its good @Stormtrooper

OpenStudy (anonymous):

I think you forgot about the subtraction.

OpenStudy (anonymous):

what do you you mean?

OpenStudy (anonymous):

(1/4x - 8) unless that's a negative 8

OpenStudy (anonymous):

how did you get 36 @kevindragonlord

OpenStudy (anonymous):

my answer was more of a guess. But if you replace the x for 36 you get 19 (But in my equation I subtract 8 from 1/4x).

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

is there a negative number in 1/4x-8+2/4x=19 ?

OpenStudy (anonymous):

idk